OpenStudy (anonymous):

use substitute to solve x=3y and 3x-5y=12

OpenStudy (anonymous):

They said to use substitute method So its simple put x = 3y in 3x - 5y = 12 3(3y) - 5y = 12 Simplify this for y value and then substitute y value in x :)

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

Substitute the equation that you will use by the value of the variable 3x-5y=12 3(3y)-5y=12 9y-5y=12 4y=12 4y/4=12/4 y=3 Then substitute the value of x x=3y x=3(3) x=9 Then check it: x=3y 9=3(3) 9=9

OpenStudy (anonymous):

Ok just ignore what I said, I did not fully read the question.

OpenStudy (anonymous):

lol ok